A diffraction grating has 3 x 106 lines per meter. The grating is illuminated by monochromatic plane waves of wavelength 600 nm at normal incidence that forms an interference pattern on a screen located 0.50 m away

a) Find the angle the first bright fringe.

b) Find the angle corresponding to the second bright fringe.

c) Find the angle of the third bright fringe.
Answer #1

slit width on grating (d) = 1/(3*!0^6) m

The equation for interfrence pattern dsin\theta = m\lambda

a) For first bright fringe

dsin\theta1 = 1\lambda

sin\theta1 = 600*10^-9 /{1/3*10^6} = 1.8

Not a possible solution

b) for Second bright fringe

d sin\theta2 = 2 \lambda

sin\theta2 = 2* 600*10^-9 /{1/3*10^6} = 3.6

also not a possible solution

c)

for Second bright fringe

d sin\theta3= 2 \lambda

sin\theta3= 2* 600*10^-9 /{1/3*10^6} = 5.4

also not a possible solution.

For this grating interefence pattern does not form with this grating.